OpenStudy (anonymous):

Four-sevenths of the children in a class earned A's on the last history test. 21 children did not earn an A. How many children earned A's? A) 49 B)77 C)3 D)28 PLEASE HELP!

OpenStudy (jack1):

3/7 = 21 so 7/7 =...?

OpenStudy (jack1):

or 1/7 = then from that, find 4/7

OpenStudy (anonymous):

I dont know it did not give me enough info.

OpenStudy (jack1):

ok so 21 = 3/7 so 1/7 = 3/7 divided by 3 so 1/7 = 21/3 = 7 now 4 / 7 got a's so 1/7 = 7 students so how many students in 4/7?

OpenStudy (anonymous):

28?

OpenStudy (jack1):

spot on champ

OpenStudy (anonymous):

WOOT ! Thank you so much :)

OpenStudy (jack1):

welcome